# The radius of statistical efficiency

### v1

@@ 4 Generalized linear models and (quasi) maximum likelihood estimation
Theorem 4.2: (RSE for QMLE) Consider a QMLE problem (22) and let $\mathsf{D}_{x}\in\mathcal{Q}$ be the distribution of $x$ with covariance $\Sigma:=\operatorname{\mathbb{E}}[xx^{\top}]$. Then, the estimate holds: $\operatorname{RSE}(\mathsf{D})=\sqrt{\lambda_{\min}\left(\Sigma|_{{\mathcal{T}}}\right)},$ (23) In particular, if for some $c_{\texttt{lb}},c_{\texttt{ub}}>0$ the inequality $c_{\texttt{lb}}\leq h^{\prime\prime}(\langle x,\beta_{\star}\rangle)\leq c_{\texttt{ub}}$ holds for $\mathsf{D}_{x}$-almost every $x$, then we have $c_{\texttt{lb}}\leq(\operatorname{RSE}(\mathsf{D}))^{2}\cdot{\rm REG}(\mathsf{D})\leq c_{\texttt{ub}}.$
Thus we see that under mild conditions, the hardness of the problem is inversely proportional to the square of the distance to the nearest ill-posed problem, ${\rm REG}(\mathsf{D})\propto\operatorname{RSE}(\mathsf{D})^{-2}$. Note that this scaling is different from the one exhibited by PCA in the previous section, ${\rm REG}(\mathsf{D})\propto\operatorname{RSE}(\mathsf{D})^{-1}$.
@@
Aside from the examples in Table 1, an interesting problem instance occurs in sparse recovery. Namely, let $\mathcal{M}$ be the submanifold of $\mathbf{R}^{d}$ comprised of $k$-sparse vectors, i.e., those which have precisely $k$ nonzero components. Then the tangent space of $\mathcal{M}$ at ${\beta}_{\star}$ is the $k$-dimensional subspace of $\mathbf{R}^{d}$ in which ${\beta}_{\star}$ is supported:
${\mathcal{T}}={\operatorname{span}}\{{e}_{i}\mid\langle{e}_{i},{\beta}_{\star}\rangle\neq 0\big{\}},$
@@
where $\{{e}_{1},\ldots,{e}_{d}\}$ denotes the standard basis of $\mathbf{R}^{d}$. Any regression problem from Table 1 constrained to $\mathcal{M}$ is ill-conditioned precisely when $\operatorname{supp}(x)\subset{v}^{\perp}$ for some ${v}\in{\mathcal{T}}$. The right-side of (23) is then the square root of the minimum eigenvalue of the submatrix of $\Sigma_{\mathsf{D}}$ whose columns and rows are indexed by the nonzero coordinates of ${\beta}_{\star}$.
@@ 5 Rank-one matrix regression problems
In this section, we compute the RSE for a number of tasks in low-rank matrix recovery—a large class of problems with numerous applications in control, system identification, recommendation systems, and machine learning. See for example [24, 17, 14, 70] for an overview. Setting the stage, consider the measurement model
$y=\langle X,M_{\star}\rangle+\varepsilon$ (24)
where the matrix $X\in\mathbf{R}^{d_{1}\times d_{2}}$ is drawn from some distribution $\mathsf{D}_{x}$, the noise $\varepsilon$ is independent of $X$ and is mean zero with variance $\sigma^{2}$, and $M_{\star}$ is a rank $r$ matrix. The goal of low-rank matrix recovery is to estimate the latent parameter $M_{\star}$ for finitely many i.i.d samples $(X_{1},y_{1}),\dots,(X_{n},y_{n})$. For the rest of the section, we focus on the simplest case of rank $r=1$ matrix recovery.
@@ 5.1 Phase retrieval
The problem of phase retrieval corresponds to (24), where the ground truth $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$ and the data $X=xx^{\top}$ matrices are symmetric and have rank one. We will let $\mathsf{D}_{x}\in\mathcal{P}_{4}(\mathbf{R}^{d})$ denote the distribution of $x\in\mathbf{R}^{d}$ and let $\mathsf{D}$ denote the joint distribution of $(x,y)$. There are two standard ways to write the phase retrieval problem as a problem of stochastic optimization. The first is simply to minimize the mean square error over rank one matrices:
$\min_{{M\succeq 0:\,\operatorname{rank}M=1}}\,f(M):=\tfrac{1}{2}\mathop{\operatorname{\mathbb{E}}}_{x,y\sim\mathsf{D}}\big{(}x^{\top}Mx-y\big{)}^{2}.$ (25)
Alternatively, one may parameterize rank one matrices as $M=\beta\beta^{\top}$ and then minimize the mean square error over the factors:
$\min_{{\beta\in\mathbf{R}^{d}}}\,f(\beta):=\tfrac{1}{8}\mathop{\operatorname{\mathbb{E}}}_{x,y\sim\mathsf{D}}\big{(}\langle x,\beta\rangle^{2}-y\big{)}^{2}.$ (26)
From the viewpoint of RSE there is no significant distinction between these two formulations. The proof of the next lemma appears in Appendix E.1.
@@
$\mathcal{E}=\bigcup_{v\in\mathbb{S}^{d-1}}\Big{\{}\mu\in\mathcal{P}_{4}(\mathbf{R}^{d}):\operatorname{supp}(\mu)\subset{\beta}_{\star}^{\perp}\cup v^{\perp}\Big{\}}.$
Fix any measure $\mu\in\mathcal{P}_{4}\setminus\mathcal{E}$ and define the matrix $\hat{\Sigma}_{\mu}=\operatorname{\mathbb{E}}_{\mu}\langle x,\beta_{\star}\rangle^{2}xx^{\top}$. Then, the estimates hold:
$c_{\texttt{lb}}\leq{\rm REG}(\mu)\cdot\lambda_{\min}(\hat{\Sigma}_{\mu})\leq c_{\texttt{ub}},$ (27)
where $(c_{\texttt{lb}},c_{\texttt{ub}})=(2\|{\beta}_{\star}\|^{2},4\|{\beta}_{\star}\|^{2})$ for problem (25) and $(c_{\texttt{lb}},c_{\texttt{ub}})=(1,1)$ for problem (26).
It remains to estimate the expression for the distance to the exceptional set $\mathcal{E}$. The following lemma shows that minimizing the expected squared distance to $\{\beta_{\star}\}^{\perp}\cup\{u\}^{\perp}$ over $u\in\mathbb{S}^{d-1}$ yields the squared $W_{2}$ distance to $\mathcal{E}$; its proof appears in Appendix E.2.
@@
Thus, using the reasoning from Lemma 5.1 and Theorem 5.1 one easily derives the following two estimates for formulation (26):
@@
${\rm REG}(\mathsf{D})^{-1}$ $=\min_{v\in\mathbb{S}^{d-1}}\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{D}_{x}}\bigg{[}\Big{\langle}x,\beta_{\star}\Big{\rangle}^{2}\big{\langle}x,v\big{\rangle}^{2}\bigg{]}.$
A moment of thought leads one to realize that for reasonable distributions, the first equation should scale as $\sqrt{\lambda_{\min}(\Sigma)}$ while the scaling of the second is at least $\lambda_{\min}(\Sigma)$. We now verify that this is indeed the case when the base distribution is Gaussian $x\sim N(0,\Sigma)$ for some covariance matrix $\Sigma\succeq 0$. To this end, define the two functions
@@
$\big{(}1-\tfrac{2}{\pi}\big{)}\lambda_{\min}(\Sigma)\leq$ $\min_{v\in\mathbb{S}^{d-1}}h_{\Sigma}(u,v)\leq\lambda_{\min}(\Sigma).$ (28)
$\lambda_{\min}(\Sigma)\cdot\langle\Sigma u,u\rangle\leq$ $\min_{v\in\mathbb{S}^{d-1}}g_{\Sigma}(u,v)\leq 3\lambda_{\min}(\Sigma)\cdot\langle\Sigma u,u\rangle$ (29)
Combining Lemma 5.1, Theorem 5.1, and Theorem 5.3 directly yields the following estimate on RSE with Gaussian initial data.
Theorem 5.4: (RSE for phase retrieval with Gaussian data) Consider problems (25) and (26) with Gaussian data $\mathsf{D}_{x}\sim N(0,\Sigma)$. Then, the estimate hold: $\sqrt{(1-\tfrac{2}{\pi})\lambda_{\min}(\Sigma)}\leq{\rm RSE}(\mathsf{D})\leq\sqrt{\lambda_{\min}(\Sigma)}.$ In particular, the following estimates hold: $c_{\texttt{lb}}\cdot\frac{1-\tfrac{2}{\pi}}{3\langle\Sigma{\beta}_{\star},{\beta}_{\star}\rangle}\leq\operatorname{RSE}(\mathsf{D})^{2}\cdot{\rm REG}(\mathsf{D})\leq c_{\texttt{ub}}\cdot\frac{1}{\langle\Sigma{\beta}_{\star},{\beta}_{\star}\rangle}$ (30) where $(c_{\texttt{lb}},c_{\texttt{ub}})=(2\|{\beta}_{\star}\|^{2},4\|{\beta}_{\star}\|^{2})$ for problem (25) and $(c_{\texttt{lb}},c_{\texttt{ub}})=(1,1)$ for problem (26).
Interestingly, we see a different scaling between $\operatorname{RSE}(\mathsf{D})$ and ${\rm REG}(\mathsf{D})$, depending on whether ${\beta}_{\star}$ aligns with the nullspace of $\Sigma$. In the regime $\langle\Sigma{\beta}_{\star},{\beta}_{\star}\rangle\gg\lambda_{\min}(\Sigma)$, we observe the scaling ${\rm REG}(\mathsf{D})\propto\operatorname{RSE}(\mathsf{D})^{-2}$, while in the regime $\langle\Sigma{\beta}_{\star},{\beta}_{\star}\rangle\approx\lambda_{\min}(\Sigma)$, the scaling is much worse ${\rm REG}(\mathsf{D})\propto\operatorname{RSE}(\mathsf{D})^{-4}$. Thus in the latter regime, the distance to the nearest ill-posed problem has a much stronger effect on the hardness of the problem.
@@ 5.2 Bilinear sensing
The problem of bilinear sensing is an asymmetric analogue of phase retrieval, that is, the ground truth matrix $M_{\star}=\beta_{1\star}\beta_{2\star}^{\top}$ and measurement data $X=x_{1}x_{2}^{\top}$ are rank one $d_{1}\times d_{2}$ matrices, where the factors $x_{1}\sim\mathsf{D}_{x_{1}}$ and $x_{2}\sim\mathsf{D}_{x_{2}}$ are independent. The standard way to write this problem as stochastic optimization is to minimize the mean square error over rank one rectangular matrices:
$\min_{{M\in\mathbf{R}^{d_{1}\times d_{2}}:\,\operatorname{rank}M=1}}\,f(M):=\tfrac{1}{2}\mathop{\operatorname{\mathbb{E}}}_{\mathsf{D}}\big{(}x^{\top}_{1}Mx_{2}-y\big{)}^{2},$ (31)
where $\mathsf{D}$ denotes the joint distribution over $(x_{1},x_{2},y)$. Throughout, we fix as the set of allowable data distributions all product measures $\mathcal{Q}=\mathcal{P}_{4}(\mathbf{R}^{d_{1}})\times\mathcal{P}_{4}(\mathbf{R}^{d_{1}})$. We disregard the factorized formulation with $M=\beta_{1}\beta_{2}^{\top}$ because it results in a continuum of minimizers that are not tilt-stable. This technical difficulty could be circumvented by introducing an additional constraint, such as $\|\beta_{1}\|=1$; however, we do not pursue this approach to simplify the exposition.
@@
=$\mathcal{E}=\Big{\{}\mu\times\nu\in\mathcal{Q}:~{}{\rm either}~{}\operatorname{\mathbb{E}}_{\mu}xx^{\top}~{}{\rm or~{}}\operatorname{\mathbb{E}}_{\nu}xx^{\top}{\rm~{}is~{}singular}\Big{\}}.$
Moreover, for any $\mu\times\nu\in\mathcal{Q}\setminus\mathcal{E}$, with $\Sigma_{1}=\operatorname{\mathbb{E}}_{\mu}x_{1}x_{1}^{\top}$ and $\Sigma_{2}=\operatorname{\mathbb{E}}_{\nu}x_{2}x_{2}^{\top},$ we have
@@
where $\gamma_{i}=\left\langle\Sigma_{i}\frac{\beta_{i\star}}{\|\beta_{i\star}\|},\frac{\beta_{i\star}}{\|\beta_{i\star}\|}\right\rangle$ for $i=1,2$.
Thus an application of Theorem A.1 immediately yields the following expression for RSE.
Theorem 5.6: (RSE for bilinear sensing) Consider the problem (31) with $\Sigma_{1}:=\operatorname{\mathbb{E}}_{\mathsf{D}_{x_{1}}}x_{1}x_{1}^{\top}$ and $\Sigma_{2}:=\operatorname{\mathbb{E}}_{\mathsf{D}_{x_{2}}}x_{2}x_{2}^{\top}$. Then it holds: ${\rm RSE}(\mathsf{D})=\min\left\{\sqrt{\lambda_{\min}(\Sigma_{1})},\sqrt{\lambda_{\min}(\Sigma_{2})}\right\}.$ (33) In particular, the following estimate holds: $\frac{\min\left\{\lambda_{\min}(\Sigma_{1}),\lambda_{\min}(\Sigma_{2})\right\}}{\min\{\gamma_{2}\lambda_{\min}(\Sigma_{1}),\gamma_{1}\lambda_{\min}(\Sigma_{2})\}}\leq\operatorname{RSE}(\mathsf{D})^{2}\cdot{\rm REG}(\mathsf{D})\leq C\cdot\frac{\min\left\{\lambda_{\min}(\Sigma_{1}),\lambda_{\min}(\Sigma_{2})\right\}}{\min\{\gamma_{2}\lambda_{\min}(\Sigma_{1}),\gamma_{1}\lambda_{\min}(\Sigma_{2})\}}$ where $C=\frac{\kappa(\Sigma_{1})\kappa(\Sigma_{2})+1}{2}$ with $\gamma_{i}=\left\langle\Sigma_{i}\frac{\beta_{i\star}}{\|\beta_{i\star}\|},\frac{\beta_{i\star}}{\|\beta_{i\star}\|}\right\rangle$ for $i=1,2$.
Similar to phase retrieval, the scaling between RSE and REG for bilinear sensing depends on the simultaneous alignment between $\beta_{i\star}$ and the least eigenvector of $\Sigma_{i}$ for $i=1,2$. In particular, when $\kappa(\Sigma_{1})\kappa(\Sigma_{2})\approx 1$ both upper and lower bounds are off by a constant and if, further, $\lambda_{\min}(\Sigma_{1})\approx\lambda_{\min}(\Sigma_{2}),$ there are two regimes: (1) when $\min\{\gamma_{1},\gamma_{2}\}\gg\lambda_{\min}(\Sigma_{1})$, then ${\rm REG}(\mathsf{D})\propto{\rm RSE}(\mathsf{D})^{-2}$, and (2) when $\min\{\gamma_{1},\gamma_{2}\}\approx\lambda_{\min}(\Sigma_{1})$, then ${\rm REG}(\mathsf{D})\propto{\rm RSE}(\mathsf{D})^{-4}.$

### v2

@@ 4 Generalized linear models and (quasi) maximum likelihood estimation
Theorem 4.2: (RSE for QMLE) Consider the QMLE problem (23) and let $\mathsf{D}_{x}\in\mathcal{Q}$ be the distribution of $x$ with second-moment matrix $\Sigma:=\operatorname{\mathbb{E}}[xx^{\top}]$. Suppose that $\mathcal{Q}$ satisfies the following closure property: for every $\mu\in\mathcal{Q}$, we have $(\mathsf{P}_{v^{\perp}})_{\#}\mu\in\mathcal{Q}$ for some minimal eigenvector $v$ of $\Sigma_{\mu}|_{{\mathcal{T}}}$. Then the identity holds: $\operatorname{RSE}(\mathsf{D})=\sqrt{\lambda_{\min}(\Sigma|_{{\mathcal{T}}})}.$ (24) In particular, if $\mathsf{D}_{x}\in\mathcal{Q}\setminus\mathcal{E}$ and for some $c_{\textup{{lb}}},c_{\textup{{ub}}}>0$ the inequality $c_{\textup{{lb}}}\leq{h^{\prime\prime}(\langle x,\beta_{\star}\rangle)^{-1}}\leq c_{\textup{{ub}}}$ holds for $\mathsf{D}_{x}$-almost every $x$, then we have $c_{\textup{{lb}}}\leq\operatorname{RSE}(\mathsf{D})^{2}\cdot{\rm REG}(\mathsf{D})\leq c_{\textup{{ub}}}.$
[…]
@@
Aside from the examples in Table 1, an interesting problem instance occurs in sparse recovery. Namely, let $\mathcal{M}$ be the submanifold of $\mathbf{R}^{d}$ comprised of $k$-sparse vectors, i.e., those that have precisely $k$ nonzero components. Then the tangent space of $\mathcal{M}$ at $\beta$$_{\star}$ is the $k$-dimensional subspace of $\mathbf{R}^{d}$ on which $\beta$$_{\star}$ is supported:
${\mathcal{T}}={\operatorname{span}}\{{e}_{i}\mid\langle{e}_{i},{\beta}_{\star}\rangle\neq 0\},$
@@
where $\{{e}_{1},\ldots,{e}_{d}\}$ denotes the standard basis of $\mathbf{R}^{d}$. The right-side of (24) in this case is the square root of the minimum eigenvalue of the submatrix of $\Sigma$ whose columns and rows are indexed by the nonzero coordinates of $\beta$$_{\star}$.
Nonsmooth variants of the QMLE problem (23) arise commonly in practice, and it is possible for the framework of this section to be extended to handle such problems by introducing more delicate regularity assumptions that guarantee the existence of an active manifold near the minimizer, as in [26]. For example, consider the LASSO least squares problem
$\min_{\beta:\,\|\beta\|_{1}\leq t}\,\tfrac{1}{2n}\|X\beta-y\|_{2}^{2},$ (25)
where $X$ is an $n\times d$ feature matrix with full column rank. Then (25) admits a unique solution $\beta$$_{\star}$, and under sufficient regularity assumptions (e.g., strict complementarity), the active manifold $\mathcal{M}$ on which (25) restricts to a smooth problem is the relative interior of the lowest-dimensional face of the $\ell_{1}$ ball $\{\beta\in\mathbf{R}^{d}\mid\|\beta\|_{1}\leq t\}$ containing $\beta$$_{\star}$, where $t=\|{\beta}_{\star}\|_{1}$. In this case, (24) applies with ${\mathcal{T}}=\operatorname{span}(\mathcal{M}-{\beta}_{\star})$.
@@ 5 Rank-one matrix regression problems
In this section, we compute the RSE for a number of tasks in low-rank matrix recovery—a large class of problems with numerous applications in control, system identification, recommendation systems, and machine learning. See, for example, [24, 17, 14, 74] for an overview. Setting the stage, consider the measurement model
$y=\langle X,M_{\star}\rangle+\varepsilon,$ (26)
where the matrix $X\in\mathbf{R}^{d_{1}\times d_{2}}$ is drawn from some distribution $\mathsf{D}_{X}$, the noise $\varepsilon$ is independent of $X$ and has zero mean and variance $\sigma^{2}$, and $M_{\star}$ is a rank-$r$ matrix. The goal of low-rank matrix recovery is to estimate the latent parameter $M_{\star}$ from finitely many i.i.d. samples $(X_{1},y_{1}),\dots,(X_{n},y_{n})$. For the rest of the section, we focus on the simplest case of rank $r=1$ matrix recovery.
@@ 5.1 Phase retrieval
The problem of phase retrieval corresponds to (26) with the ground truth $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$ and the feature data $X=xx^{\top}$ being PSD rank-one matrices. We will let $\mathsf{D}_{x}\in\mathcal{P}_{4}(\mathbf{R}^{d})$ denote the distribution of $x\in\mathbf{R}^{d}$ and let $\mathsf{D}$ denote the joint distribution of $(x,y)$. There are two standard ways to write the phase retrieval problem as a problem of stochastic optimization. The first is simply to minimize the mean squared error over PSD rank-one matrices:
$\min_{{M\in\bm{S}_{+}^{d}:\,\operatorname{rank}(M)=1}}\,f(M):=\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\mathsf{D}}\tfrac{1}{2}\big(x^{\top}Mx-y\big)^{2}.$ (27)
Alternatively, one may parameterize PSD rank-one matrices as $M=\beta\beta^{\top}$ and then minimize the mean squared error over the factors:
$\min_{{\beta\in\mathbf{R}^{d}\setminus\{0\}}}\,f(\beta):=\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\mathsf{D}}\tfrac{1}{8}\big(\langle x,\beta\rangle^{2}-y\big)^{2}.$ (28)
From the viewpoint of RSE, there is no significant distinction between these two formulations. The proof of the next lemma appears in Appendix E.1.
@@
$\mathcal{E}=\bigcup_{v\in\mathbb{S}^{d-1}}\Big\{\mu\in\mathcal{P}_{4}(\mathbf{R}^{d})\mid\operatorname{supp}(\mu)\subset{\beta}_{\star}^{\perp}\cup v^{\perp}\Big\}.$
Fix any measure $\mu\in\mathcal{P}_{4}\setminus\mathcal{E}$ and define the matrix $\hat{\Sigma}_{\mu}:={\operatorname{\mathbb{E}}}_{\mu}\big[\langle x,\beta_{\star}\rangle^{2}xx^{\top}\big]$. Then
$c_{\textup{{lb}}}\leq{\rm REG}(\mu)\cdot\lambda_{\min}(\hat{\Sigma}_{\mu})\leq c_{\textup{{ub}}},$ (29)
where $(c_{\textup{{lb}}},c_{\textup{{ub}}})=\big(\tfrac{1}{2}\|{\beta}_{\star}\|^{2}_{2},\|{\beta}_{\star}\|^{2}_{2}\big)$ for problem (27) and $(c_{\textup{{lb}}},c_{\textup{{ub}}})=(1,1)$ for problem (28).
It remains to compute the distance to the exceptional set $\mathcal{E}$. The following result shows that minimizing the expected squared distance to $\beta_{\star}^{\perp}\cup v^{\perp}$ over $v\in\mathbb{S}^{d-1}$ yields the squared $W_{2}$ distance to $\mathcal{E}$; its proof appears in Appendix E.2.
@@
In light of Lemma 5.1 and Theorem 5.1, the following equalities hold for the formulation (28):
@@
${\rm REG}(\mathsf{D})^{-1}$ $=\min_{v\in\mathbb{S}^{d-1}}\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{D}_{x}}\!\Big[\big\langle x,\beta_{\star}\big\rangle^{2}\cdot\big\langle x,v\big\rangle^{2}\Big].$
A moment of thought leads one to realize that for reasonable distributions, the first equation should scale as $\sqrt{\lambda_{\min}(\Sigma_{\mathsf{D}_{x}})}$, while the scaling of the second should be at least $\lambda_{\min}(\Sigma_{\mathsf{D}_{x}})$. We now verify that this is indeed the case when the base distribution is Gaussian with $x\sim\mathsf{N}(0,\Sigma)$ for some covariance matrix $\Sigma\succeq 0$. To this end, define the two functions
@@
$\big(1-\tfrac{2}{\pi}\big)\lambda_{\min}(\Sigma)\leq$ $\min_{v\in\mathbb{S}^{d-1}}h_{\Sigma}(u,v)\leq\lambda_{\min}(\Sigma),$ (30)
$\lambda_{\min}(\Sigma)\cdot\langle\Sigma u,u\rangle\leq$ $\min_{v\in\mathbb{S}^{d-1}}g_{\Sigma}(u,v)\leq 3\lambda_{\min}(\Sigma)\cdot\langle\Sigma u,u\rangle.$ (31)
Combining Lemma 5.1, Theorem 5.1, and Theorem 5.3 directly yields the following estimate on RSE with Gaussian feature data.
Theorem 5.4: (RSE for phase retrieval with Gaussian data) Consider problems (27) and (28) with Gaussian data $\mathsf{D}_{x}=\mathsf{N}(0,\Sigma)$. Then $\sqrt{(1-\tfrac{2}{\pi})\lambda_{\min}(\Sigma)}\leq{\rm RSE}(\mathsf{D})\leq\sqrt{\lambda_{\min}(\Sigma)}.$ Moreover, if $\mathsf{N}(0,\Sigma)\notin\mathcal{E}$, then the following estimate holds: $c_{\textup{{lb}}}\cdot\frac{1-\tfrac{2}{\pi}}{3\langle\Sigma{\beta}_{\star},{\beta}_{\star}\rangle}\leq\operatorname{RSE}(\mathsf{D})^{2}\cdot{\rm REG}(\mathsf{D})\leq c_{\textup{{ub}}}\cdot\frac{1}{\langle\Sigma{\beta}_{\star},{\beta}_{\star}\rangle}$ (32) where $(c_{\textup{{lb}}},c_{\textup{{ub}}})=\big(\tfrac{1}{2}\|{\beta}_{\star}\|^{2}_{2},\|{\beta}_{\star}\|^{2}_{2}\big)$ for problem (27) and $(c_{\textup{{lb}}},c_{\textup{{ub}}})=(1,1)$ for problem (28).
[…]
@@ 5.2 Bilinear sensing
The problem of bilinear sensing is an asymmetric analogue of phase retrieval, that is, the ground truth $M_{\star}=\beta_{1\star}\beta_{2\star}^{\top}$ and feature data $X=x_{1}x_{2}^{\top}$ are rank-one $d_{1}\times d_{2}$ matrices, where the factors $x_{1}\sim\mathsf{D}_{x_{1}}$ and $x_{2}\sim\mathsf{D}_{x_{2}}$ are independent. The standard way to write this problem as stochastic optimization is to minimize the mean squared error over rank-one rectangular matrices:
$\min_{{M\in\mathbf{R}^{d_{1}\times d_{2}}:\,\operatorname{rank}(M)=1}}\,f(M):=\mathop{\operatorname{\mathbb{E}}}_{(x_{1},x_{2},y)\sim\mathsf{D}}\tfrac{1}{2}\big(x^{\top}_{1}Mx_{2}-y\big)^{2}.$ (33)
Throughout, we fix $\mathcal{Q}=\mathcal{P}_{4}(\mathbf{R}^{d_{1}})\times\mathcal{P}_{4}(\mathbf{R}^{d_{2}})$ as the set of allowable feature data distributions. We disregard the factorized formulation with $M=\beta_{1}\beta_{2}^{\top}$ because it results in a continuum of minimizers that are not tilt-stable. This technical difficulty could be circumvented by introducing an additional constraint, such as $\|\beta_{1}\|=1$; however, we do not pursue this approach to simplify the exposition.
@@
$\mathcal{E}=\Big\{\mu\times\nu\in\mathcal{Q}\mid{\rm either}~\operatorname{\mathbb{E}}_{\mu}[xx^{\top}]~{\rm or~}\operatorname{\mathbb{E}}_{\nu}[xx^{\top}]{\rm~is~singular}\Big\}.$
Moreover, for any $\mu\times\nu\in\mathcal{Q}$ with $\Sigma_{1}:=\operatorname{\mathbb{E}}_{\mu}[x_{1}x_{1}^{\top}]$ and $\Sigma_{2}:=\operatorname{\mathbb{E}}_{\nu}[x_{2}x_{2}^{\top}]$, we have
@@
where $\gamma_{i}:=\left\langle\Sigma_{i}\frac{\beta_{i\star}}{\|\beta_{i\star}\|_{2}},\frac{\beta_{i\star}}{\|\beta_{i\star}\|_{2}}\right\rangle$ for $i=1,2$.
Thus, an application of Theorem A.1 immediately yields the following expression for RSE.
Theorem 5.6: (RSE for bilinear sensing) Consider the problem (33) with $\Sigma_{1}:=\operatorname{\mathbb{E}}_{\mathsf{D}_{x_{1}}}[x_{1}x_{1}^{\top}]$ and $\Sigma_{2}:=\operatorname{\mathbb{E}}_{\mathsf{D}_{x_{2}}}[x_{2}x_{2}^{\top}]$. Then ${\rm RSE}(\mathsf{D})={\min}\bigg\{\sqrt{\lambda_{\min}(\Sigma_{1})},\sqrt{\lambda_{\min}(\Sigma_{2})}\bigg\}.$ (35) In particular, if $\mathsf{D}_{x_{1}}\times\mathsf{D}_{x_{2}}\in\mathcal{Q}\setminus\mathcal{E}$, then the following estimate holds: $\frac{\min\{\lambda_{\min}(\Sigma_{1}),\lambda_{\min}(\Sigma_{2})\}}{\min\{\gamma_{2}\lambda_{\min}(\Sigma_{1}),\gamma_{1}\lambda_{\min}(\Sigma_{2})\}}\leq\operatorname{RSE}(\mathsf{D})^{2}\cdot{\rm REG}(\mathsf{D})\leq C\cdot\frac{\min\{\lambda_{\min}(\Sigma_{1}),\lambda_{\min}(\Sigma_{2})\}}{\min\{\gamma_{2}\lambda_{\min}(\Sigma_{1}),\gamma_{1}\lambda_{\min}(\Sigma_{2})\}}$ where $\gamma_{i}:=\left\langle\Sigma_{i}\frac{\beta_{i\star}}{\|\beta_{i\star}\|_{2}},\frac{\beta_{i\star}}{\|\beta_{i\star}\|_{2}}\right\rangle$ for $i=1,2$ and $C:=\frac{\kappa(\Sigma_{1})\kappa(\Sigma_{2})+1}{2}$.
[…]